A company's interest rate for acquiring outside capital is 6.5% compounded annually. If $35,000 must be borrowed for 5 years, what is the total amount of interest?
Answer #1

Amount = P(1+R/100)^N

= 35,000(1+6.5/100)^5

= 47953

Interest = 47953 - 35,000

= 12,953